OpenStudy (help_people):

The following function represents the value of a house after x years: f(x) = 242,000(0.96)x What does 0.96 represent? The present value of the house The value of the house after x years The increase in the value of the house per year, which is 4% The decrease in the value of the house per year, which is 4%

OpenStudy (help_people):

@Nnesha

OpenStudy (anonymous):

.96\[\frac{ 96 }{ 100 }=\left( 1-\frac{ 4 }{ 100 } \right)\] now what can you say?

OpenStudy (help_people):

i do not know? it decreases or equals @surjithayer

OpenStudy (anonymous):

it decreases by 4%

OpenStudy (help_people):

thank you XOXO may you help me w. a couple more?

OpenStudy (help_people):

@surjithayer or more likely make sure that my work is correct?

OpenStudy (anonymous):

i will try my best.

OpenStudy (help_people):

kk hold on